OpenStudy (lena772):

Is this right? Part 1: Create a scenario for an arithmetic sequence. For example, Jasmine practices the piano for __5___ minutes on Monday. Every day she _increases her practice time by ___5___. If she continues this pattern, how many minutes will she practice on the 7th day? Be sure to fill in the blanks with the words that will create an arithmetic sequence. Use your scenario to write the function for the 7th term in your sequence using sequence notation. f(1)=5 f(n)=f(1)+5n when n>1

OpenStudy (lena772):

@phi

OpenStudy (lena772):

@jdoe0001

OpenStudy (lena772):

@ranga

OpenStudy (ranga):

The nth term of an arithmetic sequence \[\Large a_n = a_1 + (n-1)d\]

OpenStudy (ranga):

Just to make the numbers different I would choose say she practices 10 minutes on Monday and increases her practice time by 5 minutes every day. The equation would become: \[\Large a_n = 10 + (n-1)*5 = 10 + 5n - 5 = 5n + 5\]

OpenStudy (ranga):

To find how much time she spends in practice on the 7th day, put n = 7 in the above formula and calculate the 7th term.

OpenStudy (lena772):

So avn=5n+5

OpenStudy (ranga):

On the nth day she will practice 5n + 5. This assumes she starts out with 10 minutes of practice on the first day and increases her practice by 5 minutes each day.

OpenStudy (lena772):

av7=5*7+5 av7=40

OpenStudy (ranga):

Yes, 40 minutes on the 7th day. And you can verify this by manually adding the minutes: Day: 1, 2, 3, 4, 5, 6, 7 Minutes 10, 15, 20, 25, 30, 35, 40 Practiced